# Conserved Quantities in Expanding Gödel Cosmology

### v1

@@ Abstract
At linear order we study perturbations to a Gödel background spacetime which includes expansion in addition to rotation. We investigate the transformation behaviour of these perturbations under gauge transformations and construct gauge invariant quantities. Using the perturbed energy conservation equation we find that there are conserved quantities in Expanding Gödel (EG) Cosmology, in particular a spatial metric trace perturbation, ${\zeta}_{\rm{SMTP}}$, which is conserved on large scales for pressureless dust. We intend to extend our discussion to a perfect fluid matter content with a view to also obtaining conserved quantities in this context.
@@ I Introduction
The earliest and perhaps best known of these cosmologies is that proposed by Kurt Gödel in 1949  GODEL1949 , which famously gave exact solutions to the Einstein Field Equations. While this cosmology includes rotation it does not allow for expansion of the universe as supported by current observations e.g.  Hubble . In 2001 Yuri N. Obukhov constructed an extension to Gödel’s cosmology which does allow for an expanding as well as rotating universe  REFEG , in which the author states it is an exact solution as per the original Gödel cosmology. Hereafter this extension is referred to as Expanding Gödel (EG) cosmology. Indeed,  HUBBLETENSION specifically cites  REFEG as the model they are likely to use in further studies in the context of rotating cosmologies.
@@
This paper is structured as follows. In the following section II we present the general governing equations under General Relativity independent of a given spacetime. In Section III we review Expanding Gödel cosmology at the background level for a pressureless dust matter content. We then extend this by adding perturbations to the Expanding Gödel background. We also examine the transformation behaviour of matter and metric variables and then construct gauge invariant quantities. In the last part of this section we look at the evolution of the spatial metric trace perturbation, ${\zeta}_{\rm{SMTP}}$. Finally in Section IV we discuss the implications of our work for studies of Expanding Gödel spacetime, the evolution of structure therein, as well as other areas of further research in Cosmology more widely in which this work may prove useful.
@@ III.5 Evolution of ${\zeta}_{\rm{SMTP}}$
IV Discussion and conclusion
@@
We derived the evolution equation for ${\zeta}_{\rm{SMTP}}$ and found that it is conserved on large scales for barotropic fluids (when $\delta P_{\mathrm{nad}}=0$). This is similar to the standard FRW result, where the similar gauge-invariant quantity, $\zeta$, is only conserved on large scales.
@@
A rotating spacetime such as EG could also contribute to the anomalous CMB dipole (see e.g. (DipoleA, ; DipoleB, )), particularly if we are situated a significant distance from the axis of rotation. Even given the magnitude of rotation has reached negligible levels in later times - and the cosmology currently conforms to FRW - this rotation in earlier epochs, and the location of its axis, could still have had a detectable effect on the CMB radiation.
Finally, recent detections by the JWST of large galaxies in the very early universe (see e.g. EarlygalaxyA ; EarlygalaxyB ) seem to disagree with current models of galaxy formation and the time-line for such formation. If in the very early universe the rotation still had significant dynamical effects on galaxy formation, as suggested by the galaxy rotation alignment results JWST , perhaps there were further effects on the evolution of these early galaxies. Possibly with the entire spacetime rotating the entire universe shared some attributes - though on a far larger scale - with accretion disks around stars, possibly compressing the distribution of matter overall in $z$-axis i.e. the axis of rotation as defined in the EG model REFEG . This compression could aid and accelerate the formation of galaxies analogous to the formation of planets within stellar accretion disks. Again, this could prove a valuable avenue of further research.
@@ Appendix B Additional Material for Perturbed EG Spacetime
$\delta g^{\mu\nu}=\begin{pmatrix}2\Phi&\frac{B^{x}}{a}&-\frac{B^{y}}{\sqrt{\sigma}X}&\frac{B^{z}}{a}\\ \\ \frac{B^{x}}{a}&-2\frac{C^{xx}}{a^{2}}&-\frac{C^{xy}}{aX}&-\frac{C^{xz}}{a^{2}}\\ \\ -\frac{B^{y}}{\sqrt{\sigma}X}&-\frac{C^{xy}}{aX}&-2\frac{C^{yy}}{\kappa X^{2}}&-\frac{C^{yz}}{aX}\\ \\ \frac{B^{z}}{a}&-\frac{C^{xz}}{a^{2}}&-\frac{C^{yz}}{aX}&-2\frac{C^{zz}}{a^{2}}\end{pmatrix}\,$. (78)

### v3

@@ Abstract
At linear order we study perturbations to a Gödel background spacetime which includes expansion in addition to rotation. We investigate the transformation behaviour of these perturbations under gauge transformations and construct gauge invariant quantities. Using the perturbed energy conservation equation we find that there are conserved quantities in Expanding Gödel (EG) Cosmology, in particular a spatial metric trace perturbation, ${\zeta}_{\rm{SMTP}}$, which is conserved on large scales for pressureless dust. We finally extend our discussion to a perfect fluid matter content to also obtain conserved quantities in this context.
@@ I Introduction
The earliest and perhaps best known of these cosmologies is that proposed by Kurt Gödel in 1949  GODEL1949 , which famously gave exact solutions to the Einstein Field Equations. While this cosmology includes rotation it does not allow for expansion of the universe as supported by current observations e.g.  Hubble . In 2000 Yuri N. Obukhov constructed an extension to Gödel’s cosmology which does allow for an expanding as well as rotating universe  REFEG , in which the author states it is an exact solution as per the original Gödel cosmology. Hereafter this extension is referred to as Expanding Gödel (EG) cosmology. Indeed,  HUBBLETENSION specifically cites  REFEG as the model they are likely to use in further studies in the context of rotating cosmologies.
@@
This paper is structured as follows. In the following section II we present the general governing equations under General Relativity independent of a given spacetime. In Section III we review Expanding Gödel cosmology at the background level for a pressureless dust matter content. We then extend this by adding perturbations to the Expanding Gödel background. We also examine the transformation behaviour of matter and metric variables and then construct gauge invariant quantities. In the last part of this section we look at the evolution of the spatial metric trace perturbation, ${\zeta}_{\rm{SMTP}}$. Finally in Section V we discuss the implications of our work for studies of Expanding Gödel spacetime, the evolution of structure therein, as well as other areas of further research in Cosmology more widely in which this work may prove useful.
@@ III.5 Evolution of ${\zeta}_{\rm{SMTP}}$
IV A Perfect Fluid EG Spacetime
Here we make a final extension to our work in EG cosmology to include a perfect fluid matter content. This could be useful when trying to examine the behaviour of perturbations - and thereby structure formation - in the early universe, when background pressure becomes significant.
IV.1 Perfect Fluid Background
The stress-energy-momentum tensor for a perfect fluid, Eq. (2), gives rise to the background energy conservation equation,
$\dot{\rho}+2H\rho+3H_{X}\rho+2HP+H_{X}P+\frac{2\kappa}{\sigma}H_{X}P=0\,.$ (70)
IV.2 Perfect Fluid Perturbations
As with the earlier pressureless dust model, III.2, split quantities into a $t$ and $x$ dependent background part where necessary, and a perturbation depending on all four coordinates. We therefore decompose the pressure $P$ as follows,
$P=\bar{P}(t,x)+\delta P(x^{\mu})\,.$ (71)
The perturbed energy conservation equation is,
[…]
IV.3 Gauge-invariant Quantities in Perfect Fluid EG Spacetime
Our gauge-invariant spatial metric trace perturbation on co-moving, uniform density hypersurfaces, ${\zeta}_{\rm{SMTP}}$, is constructed as in III.4.1 giving Eq. (62) as before. Differences only become apparent when examining the evolution of ${\zeta}_{\rm{SMTP}}$, due to the form of the perturbed energy conservation equation, Eq. (LABEL:PertEconsPerfectEG), which now contains additional terms including the background pressure. Also, since the density perturbation on uniform spatial metric trace perturbation hypersurfaces is derived simply from the transformation behaviour of the density perturbation to give Eq. (64) the relation between it and ${\zeta}_{\rm{SMTP}}$ remains Eq. (LABEL:GIdenspertZeta2).
IV.4 Evolution of ${\zeta}_{\rm{SMTP}}$ in a Perfect Fluid
If we take $\delta P=\delta P_{\mathrm{nad}}$ on uniform density hypersurfaces and use the large scale limit the evolution equation for ${\zeta}_{\rm{SMTP}}$, is found using the time derivative of Eq. (62), the perturbed energy conservation equation for a perfect fluid, Eq. (LABEL:PertEconsPerfectEG), and the background conservation equation for a perfect fluid, Eq. (70), as
${\dot{\zeta}}_{\rm{SMTP}}=\frac{1}{3}\left(2H+H_{X}+\frac{2\kappa}{\sigma}H_{X}\right)\frac{\delta P_{\mathrm{nad}}}{\bar{\rho}+\bar{P}}\,.$ (73)
which is conserved for barotropic fluids as with Eq. (69).
We do not extend the work here to include a cosmological constant in our matter content since Dark Energy domination only occurs at relatively late times. By the epoch of Dark Energy domination it is likely rotation will have reduced to negligible levels and the spacetime will have effectively reduced to FRW. The evolution of perturbations and structure formation with Dark Energy has already been extensively researched using this cosmology.
V Discussion and conclusion
@@
We derived the evolution equations for ${\zeta}_{\rm{SMTP}}$ and found that it is conserved on large scales for barotropic fluids (when $\delta P_{\mathrm{nad}}=0$). This is similar to the standard FRW result, where the similar gauge-invariant quantity, $\zeta$, is only conserved on large scales.
@@
A rotating spacetime such as EG could also contribute to the anomalous CMB dipole (see e.g. (DipoleA, ; DipoleB, )), particularly if we are situated a significant distance from the axis of rotation. Even given the magnitude of rotation has reached negligible levels in later times - and the cosmology currently conforms to FRW - this rotation in earlier epochs, and the location of its axis, could still have had a detectable effect on the CMB radiation. For an interesting review of some of the observations which are in tension with standard FRW cosmology see e.g. Alurietal .
Finally, recent detections by the JWST of large galaxies in the very early universe (see e.g. EarlygalaxyA ; EarlygalaxyB ) seem to disagree with current models of galaxy formation and the time-line for such formation. If in the very early universe the rotation still had significant dynamical effects on galaxy formation, as suggested by the galaxy rotation alignment results JWST , perhaps there were further effects on the evolution of these early galaxies. Possibly with the entire spacetime rotating the entire universe shared some attributes - though on a far larger scale - with accretion disks around stars, possibly compressing the distribution of matter overall in the $z$-axis i.e. the axis of rotation as defined in the EG model REFEG . This compression could aid and accelerate the formation of galaxies analogous to the formation of planets within stellar accretion disks. Again, this could prove a valuable avenue of further research.
@@ Appendix B Additional Material for Perturbed EG Spacetime
$\delta g^{\mu\nu}=\begin{pmatrix}2\Phi&\frac{B_{x}}{a}&-\frac{B_{y}}{\sqrt{\sigma}X}&\frac{B_{z}}{a}\\ \\ \frac{B_{x}}{a}&-2\frac{C_{xx}}{a^{2}}&-\frac{C_{xy}}{aX}&-\frac{C_{xz}}{a^{2}}\\ \\ -\frac{B_{y}}{\sqrt{\sigma}X}&-\frac{C_{xy}}{aX}&-2\frac{C_{yy}}{\kappa X^{2}}&-\frac{C_{yz}}{aX}\\ \\ \frac{B_{z}}{a}&-\frac{C_{xz}}{a^{2}}&-\frac{C_{yz}}{aX}&-2\frac{C_{zz}}{a^{2}}\end{pmatrix}\,$. (82)